OpenStudy (anonymous):

In a right triangle ABC, the legs are AB = 5 and BC = 12, and in the right triangle DEF, the hypotenuse is DE = 10, and EF = 8. Which is DF + AC

OpenStudy (anonymous):

\[5^2+12^2=25+144=169\] so hypotenuse of first triangle is \[13\] you will often see a 5-12-13 right triangle.

OpenStudy (anonymous):

second triangle is a 3-4-5 right triangle, only in this case you have 6-8-10

OpenStudy (anonymous):

so i guess you add to get 6+13=19

OpenStudy (anonymous):

AC is hypotenuse in the triangle ABC \[AC=\sqrt{AB ^{2}+BC^{2}}\] - Pythagorean theorem \[AC=\sqrt{5 ^{2}+12^{2}}\] \[AC=\sqrt{25+144}\] \[AC=\sqrt{169}\] \[AC=13\] In DEF triangle DE=10 is hypotenuse and leg EF =8 \[DE ^{2}=EF ^{2}+DF ^{2}\] \[DF ^{2}=DE ^{2}-EF ^{2}\] \[DF=\sqrt{DE ^{2}-EF ^{2}}\] \[DF=\sqrt{10^{2}-8^{2}}\] \[DF=\sqrt{100-64}\] \[DF=\sqrt{36}\] DF=6 DF+AC=6+13=19

OpenStudy (anonymous):

can any of you help me on my next question please